Assuming that a firm in a perfect competitive industry has the following total costs schedule:
Output Total Cost
10 110
15 150
20 180
25 225
30 300
35 385
40 480


a. Calculate a marginal cost and an average cost schedule for the firm
b. If the prevailing market price is $17 per unit, how many units will be produced and sold? What are profits per unit? What are total profits?
c. Is the industry in long0run equilibrium at this price?
1
Expert's answer
2014-09-05T12:27:28-0400
a. Calculate a marginal cost and an average cost schedule for the firm
ATC = TC/Q, MC = deltaTC/deltaQ
Q TC MC ATC
10 110 - -
15 150 8 10
20 180 6 9
25 225 9 9
30 300 15 10
35 385 17 11
40 480 19 12
b. If the prevailing market price is $17 per unit, 35 units will be produced and sold, because in equilibrium MC = MR = P = 17. Profits per unit are: AP = P - ATC = 17 - 11 = $6. Total profits are TP = (P - ATC)*Q = (17 - 11)*35 = $210.
c. The industry is not in long-run equilibrium at this price, because the firms earn profits. And in the long-run they will earn zero profits.